### v1

# MnBr2 on the graphene on Ir(110) substrate: growth, structure, and super-moiré

Abstract
Single-layer MnBr2 is grown on graphene (Gr) supported by Ir(110) and investigated using low-energy electron diffraction, scanning tunneling microscopy, and spectroscopy. The structure and epitaxial relationship with the substrate are systematically characterized. The growth morphology strongly depends on the growth temperature, evolving from fractal to dendritic and eventually to compact dendritic–skeletal islands, reflecting changes in the underlying surface diffusion processes. MnBr2 on Gr/Ir(110) constitutes a three-lattice system, giving rise to a super-moiré pattern –a moiré of moirés. Due to the involvement of lattices of differing symmetries and the partial electronic transparency of Gr, a "virtual" moiré formed by MnBr2 and Ir(110) contributes to the super-moiré formation. Ab initio calculations play a crucial role in understanding the complexity of super-moiré. Moreover, the pronounced variation in the apparent height with tunneling conditions for the magnetic insulator is explained based on the measured electronic structure.
I Introduction
Transition metal dihalides (TMDHs) are a class of materials that often exhibit magnetic ordering in their single-layer form owing to their unique electronic structures and strong spin-orbit coupling [1, 2, 3, 4, 5, 6]. Since the experimental discovery of intrinsic magnetic ordering in two-dimensional materials, single-layer TMDHs have attracted substantial research interest due to their potential as low-dimensional magnets and their applicability in spintronics and quantum technologies [7, 8, 9].
Beyond their intriguing magnetic properties, epitaxially grown single-layer TMDHs on graphene (Gr) undergo spontaneous self-trapping of charges, forming quasiparticles with extremely long lifetimes, identified as polarons [10, 11, 12, 13]. The formation of these quasiparticles is of fundamental and technological significance because they can mediate not only charge but also spin information in magnetic materials. However, the exact nature of these quasiparticles remains debatable because of their exceptionally long lifetimes. Some studies attribute them to small polarons [10, 11, 13, 14], whereas others explain them as solitons [15, 16]. Because these quasiparticles are observed in epitaxially grown single-layers TMDHs, understanding their exact nature requires a detailed analysis of the TMDH-Gr epitaxial relationship.
The investigation of single-layer TMDHs is often limited to first-principles calculations, with few experimental validations. An example lacking experimental realization is 2D manganese dibromide (MnBr2). For bulk MnBr2 a double-row striped antiferromagnetic ordering in the layers below 2.15 K was measured. This order becomes sinusoidally modulated between 2.15 K and 2.35 K before turning paramagnetic [17, 18, 19, 20]. Bulk MnBr2 crystallizes in a layered trigonal CdI2-type structure (space group: $\rm{P\bar{3}m1}$), where the layers are separated by van der Waals (vdW) gaps. In each layer, Mn2+ ions occupy an octahedral site coordinated by six Br- ions in an edge-sharing arrangement. Above 623 K bulk MnBr2 undergoes a phase transition in the layer stacking sequence, transforming the trigonal CdI2-type to rhombohedral CdCl2-type structure (space group: $\rm{R\bar{3}m}$) [21]. Due to the Mn2+ high-spin d5 configuration in dihalides, Mn possesses the highest magnetic moment of $\approx 5\mu_{\mathrm{B}}$ among the TMDHs. The combination of the low electronegativity of Mn (enhancing ionic interactions with halides) and its half-filled 3d5 configuration (suppressing electron hopping) creates favorable conditions for charge localization and enhances electron-phonon coupling. These characteristics suggest that polarons could also form within Mn dihalides.
[…]
The preparation of single-layer MnBr2 has not yet been realized. Preparation by exfoliation may be challenging because the strength of the in-plane bonds of TMDHs is weaker than that of materials such as Gr or transition metal dichalcogenides. Presumably for this reason, most experimental studies on single layers of TMDHs have been conducted using samples grown by molecular beam epitaxy (MBE) [26, 27, 28, 29, 30, 11, 10, 31, 32, 33, 12, 34, 35, 36]. Therefore, MBE of MnBr2 is also the method of choice used in this study.
Here, we present a comprehensive characterization of single-layer MnBr2, providing a foundation for future studies of its magnetic and polaronic properties. The investigation includes the determination of the structure and layer stacking, the analysis of growth or annealing temperature-dependent morphology linked to the underlying nucleation and diffusion processes, a systematic investigation of tunneling parameter-dependent layer height and its link to the electronic structure, and lastly, the determination of the super-moiré formed by the MnBr2 layer with the substrate moiré.
In recent years, high-order moiré patterns that emerge from the geometrical interference of more than two lattices, so-called super-moiré or moiré of moiré, have captured significant research interest within the condensed matter physics community because they provide an exceptional platform for realizing exotic quantum states [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. For instance, super-moiré systems can host topological flat bands [49, 50, 51, 47, 48] that enhance the electron-electron interactions, leading to the emergence of strongly correlated states, such as unconventional superconductors [52], fractional Chern insulators [53], quantum spin liquids [54], and Wigner crystals [55]. In $k$-space, a moiré results from an integer linear combination of vectors $\vec{k}[\mathrm{A}]$ and $\vec{k}[\mathrm{B}]$ belonging to two different lattices A and B, resulting in moiré vectors $\vec{k}[\mathrm{A-B}]$. Consequently, a super-moiré is considered an integer linear combination of two different sets of moiré vectors that originate from two different moirés [41, 42]. This requires a minimal set of three lattices, A, B, and C, with sets $\vec{k}[\mathrm{A-B}]$ and $\vec{k}[\mathrm{B-C}]$. To date, super-moiré systems have been mostly limited to twisted triple-layer Gr [45, 46, 47, 40] or heterostructures between Gr and hexagonal boron nitride [41, 42, 43, 44, 38], where all the lattices have hexagonal symmetries.
Here, we investigate the case of MnBr2/Gr/Ir(110), in which two lattices are hexagonal (Gr and MnBr2), but one is rectangular, namely the surface lattice of Ir(110). The resulting super-moiré not only exhibits intriguing complexity, but is also formed differently than outlined above. Instead of resulting from the combination of moiré vectors $\vec{k}[\mathrm{A-B}]$ and $\vec{k}[\mathrm{B-C}]$, its dominating Fourier coefficient results from a vector $\vec{k}[\mathrm{A-C}]$, that is, in the present case $\vec{k}[\mathrm{MnBr_{2}/Ir(110)}]$. This is surprising, since Ir(110) and the MnBr2 layer are not in contact with each other. Nevertheless, this study provides a tractable explanation for this phenomenon.
II Results and discussion
II.1 Structure
Figure 1: Structural characterization of MnBr2. (a) Contrast-inverted 90 eV LEED pattern of MnBr2 after the growth of 1.1 ML at 400 K on Gr/Ir(110). First-order Ir, Gr, and MnBr2 reflections are encircled in magenta, green, and orange, respectively. Reciprocal Ir, Gr and MnBr2 primitive translations are indicated. (b) STM topography corresponding to the sample in (a). (c) Atomically resolved STM image of MnBr2. (c) and (d) Top and side view ball models of the structure and bilayer stackings of MnBr2. All subfigures are oriented as indicated in (a), with the [001]-direction of Ir being horizontal. STM images are obtained at 300 K with (b) $V_{\mathrm{b}}$ = -2 V, 50 pA and (c) $V_{\mathrm{b}}$ = -2 V, 500 pA. Image sizes: (b) 180 nm $\times$ 200 nm, and (c) 7.5 nm $\times$ 7.5 nm.
The first insight into the MnBr2 structure and epitaxial relationship with Gr/Ir(110) is provided through the low-energy electron diffraction (LEED) pattern in Figure 1(a), taken after the growth of 1.1 ML at 400 K. The LEED pattern displays the first-order diffraction spots of the rectangular lattice of unreconstructed Ir(110) [encircled in magenta, with primitive reciprocal translations $\vec{a}^{*}_{Ir,1}$ and $\vec{a}^{*}_{Ir,2}$], hexagonal Gr lattice [encircled in green, with primitive reciprocal translations $\vec{a}^{*}_{Gr,1}$ and $\vec{a}^{*}_{Gr,2}$], and hexagonal MnBr2 lattice [encircled in orange, primitive reciprocal translations $\vec{a}^{*}_{MnBr_{2},1}$ and $\vec{a}^{*}_{MnBr_{2},2}$)]. Gr and Ir(110) are epitaxially aligned with a Gr zigzag direction parallel to $001]_{\mathrm{Ir}} [$56]. MnBr2 forms a hexagonal 2D lattice epitaxially aligned to Gr, with the two hexagonal lattices rotated by 30 with respect to each other. However, the epitaxy is not perfect, as manifested by the azimuthal smearing of the MnBr2 diffraction spots and the presence of a faint, barely visible diffraction ring (highlighted by the dotted circle segment and black arrow). Using the Ir diffraction spots as a reference, the analysis of multiple LEED patterns yields a MnBr2 lattice parameter of 3.90 Å $\pm$ 0.01 Å. This value is in excellent agreement with both the experimental bulk MnBr2 lattice constant of 3.873 Å [57] and the theoretically predicted values for single-layer MnBr2 [3, 25, 24]. Other spots not mentioned and visible in the LEED pattern can be traced back to the moiré of Gr on Ir(110) [56].
Figure 1(b) presents a scanning tunneling microscopy (STM) overview topograph of the same sample characterized by LEED, revealing two large first-layer MnBr2 islands partially overlaid by smaller coalesced second-layer islands on top. The islands exhibit a triangular morphology with mostly straight edges, consistent with the C3v symmetry expected for a single layer of MnBr2. At $V_{\mathrm{b}}=-2.0$ V, the apparent height of the first-layer island is measured to be 4.15 Å [see Figure S1 of the supporting information (SI)]. This value is lower than the expected apparent height of the TMDH on Gr. The underlying reason for this discrepancy is discussed in detail in the Apparent Height and Density of States section. The contrast-enhanced insets of the first-layer islands in Figure 1(b) also display different stripe patterns. The stripe patterns are the super-moirés of single-layer MnBr2 with the underlying substrate, which are discussed in more detail in the Super-moiré section. The different orientations and appearances of the two stripe patterns result from the different orientations of the MnBr2 islands. They are rotated by $\approx 6^{\circ}$ with respect to each other and thus have different orientations with respect to the substrate.
Taking a closer look at Figure 1(b), it is obvious that the first-layer MnBr2 islands point in nearly opposite directions, where the triangular tip of the island in the upper part of the image points to the left and that in the lower part points to the right. The nearly opposite orientation of the large islands is consistent with the three-fold symmetric MnBr2 growing on six-fold symmetric Gr. The fact that the orientations are not exactly opposite is consistent with the azimuthally smeared MnBr2 reflections in the LEED pattern in Figure 1(a).
Figure 1(c) shows an atomically resolved STM image of single-layer MnBr2, which confirms a hexagonal lattice that is consistent with the LEED results. In addition to the atomic-scale corrugation of $\approx$ 0.2 Å, the image exhibits nanoscale modulations manifested as narrow and wide vertical stripes. The vertical stripes are the same as those visible in the contrast-enhanced insets of Figure 1(b).
The second-layer islands in Figure 1(b) have two orientations with respect to the first-layer islands. Either their triangular envelope aligns with their base or opposes it. From the analysis of many STM topographs, it is found that aligned and opposing envelopes of second-layer islands are present in nearly equal proportions. As visualized in the ball model of Figure 1(d), the aligned envelopes correspond to bulk MnBr2 which crystallizes in the CdI2 structure [17], also denoted 1T-MnBr2. In 1T-MnBr2, layers composed of octahedrally coordinated Mn ion planes sandwiched between Br ion planes are stacked exactly on top of each other. The envelopes of opposite orientations indicate a stacking fault. Deviations from 1T stacking are well known for bulk MnBr2: above 623 K MnBr2 in the CdCl2 structure is the stable phase, that is, the layers are rhombohedrally stacked, giving rise to a tripled c-axis parameter [21]. This implies small free energy differences for variations in stacking, consistent with the formation of stacking faults under the non-equilibrium conditions of MBE growth. In agreement with our observations, DFT calculations in the BiDB database [58] found identical energies for the two experimental MnBr2 stackings within the precision of DFT, that is, an energy difference of only 0.1 meV /Å2.
Figure 2: (a)-(c) STM topographs of MnBr2 on Gr/Ir(110) after deposition of $\approx 0.6$ ML at 80 K, 300 K, and 400 K, respectively. The sample in (a) was briefly annealed to $\approx 150$ K during transfer to the STM and then imaged at 77 K. Images in (b) and (c) were acquired at 300 K. The white arrows in (b) indicate the dendritic growth directions of the primary, secondary, and ternary dendrite branches. STM imaging parameters: $V_{\mathrm{b}}=-2$ V and $I_{\mathrm{t}}=50$ pA. Image sizes: 160 nm $\times$ 160 nm.
II.2 Morphological Evolution with Temperature in Growth and Annealing
Figure 2 compares the morphology of MnBr2 as a function of growth temperature. Figure 2(a), taken at 77 K after growth at 80 K with unavoidable annealing to $\approx 150$ K in between due to sample transfer, reveals fractal island growth with smooth edges and no preferential orientation. The smooth edges suggest limited edge diffusion during the annealing period. In the absence of annealing at $\approx$ 150 K, the islands would probably have retained even more irregular fractal shape. As shown in the inset of Figure 2(a), most of the islands exhibit a single well-defined moiré orientation, except when grown together; an example of this is found in the upper left corner of the contrast-enhanced inset. The inset shows stripes of the Gr/Ir(110) substrate moiré. The average periodicity is 3.3 nm [56], although due to the magnification effect of the Gr/Ir(110) moiré for strain and defects, substantial variations occur around this average. In any case, even when the MnBr2 moiré stripes are parallel to those of the substrate, their wavelength is substantially smaller. Second-layer islands are largely absent. The average distance between islands of about 25 nm is much smaller than the typical separation of about 150 nm between steps or other surface defects. This implies the dominance of homogeneous nucleation.
Figure 2(b), acquired at 300 K after growth at 300 K, shows a segment of a large dendritic island. While the main visible branch of the dendrite grows from bottom to top [direction labeled 1 in the inserted sketch in Figure 2(b)], secondary branches grow away from it [directions labeled 2], maintaining angles of approximately 120. These angles correspond to the preferred growth directions, which are consistent with the three-fold symmetry of the crystal structure. The island edges are smooth on a much larger length scale than in Figure 2(a) and display undulations. Sharp grooves were observed between the secondary branches. Their origin can be traced back to the limited molecule supply from the diffusion field, which is preempted by the rapidly growing secondary branches and islands in the second layer. Some of the grooves are pinched, causing holes. The onset of tertiary branches is visible in Figure 2(b) [directions labeled 3]. From the dendritic appearance, the rather smooth edges, and the simultaneous existence of grooves, it must be concluded that step edge diffusion is effective on length scales of the order 5 nm, but limited on large length scales of the order 50 nm. Searching the sample by moving the STM scan frame shows dendritic growth all over, but the substrate steps and defects give rise to heterogeneous nucleation, making a quantitative assessment of the nucleation density impossible to achieve. Qualitatively, the strong reduction in the island number density with increasing growth temperature is consistent with classical nucleation theory [59]. A large number of second-layer islands nucleated on top of the dendritic island. This observation implies that the diffusion of MnBr2 molecules on the first-layer MnBr2 islands is hampered compared to diffusion on Gr/Ir(110), likely due to a larger activation barrier for molecule migration. Second-layer islands are seen to be reduced in number density next to the steps of the first-layer islands. This implies that for molecules diffusing on a first-layer island, the edge acts as a sink for molecules by incorporating them. Consequently, if there is any Ehrlich-Schwoebel barrier [60, 61, 59] for the descent of molecules from the first-layer island, it must be very small.
Figure 2(c) taken at 300 K and grown at 400 K shows two first-layer islands grown together. The right island is a good example of dendritic-skeletal growth [59]. Its envelope is triangular, but its triangular sides are concave. The diffusion field has the strongest concentration gradient at the triangle tips; thus, the tips receive the largest supply of molecules. Although diffusion is efficient, transport along the edge is not fast enough to keep the edges straight, causing concave edges. Due to the higher growth temperature, the nucleation density of second-layer MnBr2 islands is smaller compared to the 300 K case. The zone denuded of second-layer islands adjacent to the first-layer island edges has grown in width, consistent with enhanced molecule diffusion at higher temperature. The second-layer islands are compact, and if not coalesced, of hexagonal to triangular in shape, since, due to their small size, diffusion is efficient in keeping the edges straight.
Taken together, the inferred picture of island nucleation and shape evolution is consistent with a strong increase with the growth temperature of (i) molecule diffusion on Gr/Ir(110) and (ii) step edge diffusion. The much higher diffusion coefficient of molecules on Gr/Ir(110) compared to that on the first-layer MnBr2 islands implies that homogeneous nucleation has already ceased for the first-layer islands at 300 K, while for the second-layer islands, it remains homogeneous throughout the investigated temperature range. The higher molecular diffusion coefficient on Gr/Ir(110) also indicates a weaker interaction between MnBr2 and Gr than between layers of MnBr2.
Figure 3: (a)-(d) The MnBr2 sample of Figure 2(b) grown at 300 K is annealed in isochronal steps of 120 s successively to 370 K, 470K̇, 570 K, and 670 K, respectively. The white arrows in (c) highlight the two pinning centers. Imaging is conducted 300 K with $V_{\mathrm{b}}$ = -2 V, $I_{\mathrm{t}}=50$ pA, Images sizes: 160 nm $\times$ 160 nm.
Figure 3 presents an isochronal annealing sequence of the sample shown in Figure 2(b) grown at 300 K. The first two annealing steps to 370 K [Figure 3(a)] and 470 K [Figure 3(b)] do not cause significant morphological changes. The island edges tend to become smoother, and at 470 K, the deep grooves of the dendrite branches disappear. Annealing to 570 K as represented by Figure 3(c) shows significant morphological changes. The second-layer islands disappeared. First-layer island steps show bright spots connected by concave-step segments. These observations indicate the onset of MnBr2 molecule evaporation from the island edges. Due to their small radii of curvature, second-layer islands have the highest vapor pressure and disappear first. The bright spots at the first-layer island edges are pinning centers, presumably Mn clusters resulting from MnBr2 decomposition. The evaporation of MnBr2 molecules between the pinning centers causes these step segments to become concave, thereby lowering the vapor pressure at the edges. After annealing at 670 K (Figure 3(d)), no MnBr2 remains on the surface, indicating that complete evaporation to the gas phase has occurred. This result is consistent with the Knudsen cell temperature of 670 K used for the sublimation of MnBr2.
II.3 Super-moiré
Figure 4: Ball model of super-moiré formation through overlay of the Ir(110) surface lattice (starting left), the Gr lattice (starting right), and one Br sublattice of MnBr2 (starting at the bottom). The MnBr2 lattices are rotated counterclockwise in (a) by $0.5^{\circ}$ and in (b) by $6.5^{\circ}$ with respect to the Ir[11¯0][$\bar{1}$0] direction.
The first insight into the moiré pattern in the MnBr2/Gr/Ir(110) system is obtained by overlaying ball-model representations of the three constituent lattices, as shown in Figure 4. The Ir(110) surface lattice (left) overlaps in the central area of both subfigures with the Gr honeycomb lattice (right). On these lattices, one Br (or equivalently the Mn) sublattice is overlaid (green balls, lattice shifted a bit down) at two different angles. On the left of each subfigure, a clean stripe moiré pattern is observed, and on the right, a hexagonal moiré pattern evolves, as indicated by an orange double arrow and a pink rhombus, respectively. In the region where all three lattices overlap, the stripe pattern survives but becomes modulated along the stripes due to the additional MnBr2 lattice. The stripe pattern is the dominant feature in the moiré patterns highlighted in Figure 1(b) and 2(a). The orientation and width of the stripes in the ball model are well matched to those highlighted in Figure 1(b). The fit parameter used was the slight misorientation of the MnBr2 lattice with respect to Gr/Ir[110], with angles of $0.5^{\circ}$ and $6.5^{\circ}$.
Figure 5: Fast Fourier transformation (FFT) analysis of the MnBr2/Gr/Ir(110) super-moiré. (a) Moiré-resolution STM topography of MnBr2. (b) In the upper left, the FFT of the STM image shown in (a) is presented, to the right of which is an enlarged, color-coded schematic model of the FTT. The relevant $\vec{k}$ vectors are indicated. (c-e), (g), and (h) are inverse FFTs using the spots represented by (c) pink, (d) orange, (e) blue, (g) purple, and (h) green dots in (b). (f) Sum of images shown (d) and (e). STM imaging parameters in (a) are $V_{\mathrm{b}}=-1$ V and $I_{\mathrm{t}}=50$ pA. Image size is 50 nm $\times$ 50 nm.
The super-moiré patterns in the MnBr2/Gr/Ir(110) system exhibit additional complex features beyond those visualized in the ball model. To identify the components of the super-moiré, we performed a fast Fourier transform (FFT) analysis of the high-resolution topography of the super-moiré, as shown in Figure 5(a), which is taken on the island visible in the upper part of Figure 1(b). The height profile reveals a substantial corrugation of $\approx$ 0.7 Å (see Figure S2 of SI). The FFT of Figure 5(a), presented in Figure 5(b), displays multiple peaks. The central part of the FFT limited to small $\vec{k}$ or large wavelengths, is represented in the upper left corner of Figure 5(b). A color-coded, enlarged, and schematic version of the pattern is shown in Figure 5(b). The inverse FFT analysis of sets of equally colored peaks provides insight into the main components of the super-moiré. The inverse FFT is shown in Figure 5(c) when selecting the hexagonal FFT peaks (encircled in pink). The image represents the hexagonal moiré between Gr and MnBr2 already visible in the ball model of Figure 4(a) by the pink rhombus. It has a lattice parameter of 2.5 nm. The inverse FFT is shown in Figure 5(d) when the two most pronounced Fourier peaks (orange) are selected. The image represents the interference between MnBr2 and Ir(110). It has a row spacing of 2.7 nm and gives rise to a pronounced line pattern in the MnBr2/Gr/Ir(110) super-moiré. This line pattern is clearly visible in Figure 4(a) on the left, where the Ir(110) surface rows overlap with the MnBr2 lattice. The strength of this pattern in the super-moiré is remarkable considering that Ir(110) and MnBr2 are separated by the Gr layer. The origin of this phenomenon is further analyzed in this section. Selecting the blue Fourier peaks is shown in Figure 5(e). The stripes have the larger periodicity of the Gr/Ir(110) moiré, which is 3.3 nm [56]. This periodicity is also visible in between the MnBr2 islands in the contrast-enhanced inset of Figure 2(a). The superposition of the two stripe patterns results in a beating pattern with a period of 13.5 nm. Superposition can be achieved by simply adding the two images, as shown in Figure 5(f), or by selecting the green spots in the FFT to create an inverse FFT, as shown in Figure 5(g). The green spots are due to the difference in the wave vector of the blue and orange spots [or the sum of the blue and orange $\vec{k}$ vectors in Figure 5(b)]. The modulation with a period of 13.5 nm is visible in the STM topograph of Figure 5(a). Finally, when selecting the purple spots resulting from the addition of $\vec{k}[\mathrm{Gr/Ir(110}]+\vec{k}[\mathrm{MnBr_{2}/Ir(110}]$ and similar sums (or differences), the inverse FFT is shown in Figure 5(h). The beating amplitudes of both, the vertical and stripe patterns can be recognized. The horizontal stripes are phase-shifted by $\pi$ between the segments of 13.5 nm periodicity, as in the STM topograph of Figure 5(a).
Figure 6: (a) Atomically resolved moiré of Gr/Ir(110). The inset shows the height profile along the red line. (b) DFT-calculated charge density isosurface with $2.1\times 10^{-10}$ electrons/Å3 in an average height of $3.6$ Å above the surface. The grayscale encodes the variation in isosurface height. (c) Ball model of atomic arrangement in the supercell. Navy blue: bottom-layer Ir atoms; light blue balls: top-layer Ir atoms; gray balls: C atoms. A black rhombus in all subfigures indicates the supercell size in the DFT calculations. STM imaging parameters in (a) $V_{\mathrm{b}}=100$ mV and $I_{\mathrm{t}}=1$ nA. Image size is 10 nm $\times$ 8 nm.
After establishing the main elements of the super-moiré, it remains to be understood why the Fourier component corresponding to a "virtual" moiré of MnBr2 and Ir(110) is so pronounced in the super-moiré, although the two lattices are not in contact. The answer is encoded in the atomically resolved STM topography of Gr on Ir(110), as shown in Figure 6(a). The topograph clearly shows the rectangular $3.3\,\mathrm{nm}\times 1\,\mathrm{nm}$ unit cell of the Gr/Ir(110) moiré and the corresponding stripe periodicity. However, in addition to this, the Ir(110) atomic row spacing of 3.84 Å is imprinted onto the moiré. The densely packed Ir(110) rows run vertically under Gr in Figure 6(a). The imprinted periodicity is evident from the line profile in the inset of Figure 6(a), which is on a horizontal ridge along the red line in Figure 6(a) and displays a 3.84 Å periodicity. The same Ir row periodicity is also observed in the horizontal grooves. Our findings are supported by the DFT calculations of the total charge density of the Gr/Ir(110) moiré displayed in Figure 6(b). The total charge density was calculated using DFT for the supercell indicated by a black rhombus in Figure 6(b), of which the atomic structure is shown in Figure 6(c). The charge density isosurface with $2.1\times 10^{-10}$ electrons/Å3 is at an average height of 3.6 Å above the surface. Darker gray represents smaller and lighter gray larger heights. The total charge density isosurface not only resembles the STM topograph quite well but also has the Ir atomic row periodicity imprinted as strongly as in the STM topograph. In particular, the Gr lattice is visually absent in the STM topography in Figure 6(a) and the DFT-calculated charge density in Figure 6(b). In the next section, we will see that MnBr2 is a large-band-gap insulator, for which our tunneling conditions are within the bandgap. Thus, the STM signal is dominated by tunneling into the interface of the bottom Br sublattice of MnBr2 in contact and hybridized with the electronic structure of Gr/Ir(110) shown in Figure 6. In this sense, the super-moiré observed on the MnBr2 islands is an electronic super-moiré structure situated below the MnBr2 surface. The dominance of the Fourier components related to $\vec{k}[\mathrm{MnBr_{2}/Ir(110)}]$ is a consequence of the periodicity of the top layer Ir(110) rows, which are present with a large modulation in the Gr/Ir(110) moiré.
II.4 Apparent Height and Density of States
Figure 7: (a) Apparent height of single-layer MnBr2 with respect to Gr for $I_{\mathrm{t}}=20$ pA plotted as a function of $V_{\mathrm{b}}$. (b) Constant-height $\mathrm{d}I/\mathrm{d}V$ spectra of single-layer and bilayer MnBr2. The spectra are obtained with $V_{st}$ = 2.5 V, $I_{st}$ = 100 pA, $f_{mod}$ = 667 Hz, and $V_{mod}$ = 20 mV.
The apparent STM height of single-layer MnBr2 on Gr/Ir(110) is plotted in Figure 7(a) as a function of the bias voltage $V_{\mathrm{b}}$. Over a large range from -4 V to 2 V it is rather insensitive to $V_{\mathrm{b}}$ with an almost constant value of $\sim$4 Å. Only above $V_{\mathrm{b}}=2\,V$ it increases, reaches at $V_{\mathrm{b}}$ = 3.2 V a maximum of 5.95 $\pm$ 0.05 Å and then decreases again. We note that the precise apparent heights measured depend on $I_{\mathrm{t}}$ and the tip state, but its $V_{\mathrm{b}}$-dependence, as represented in Figure 7(a) is generic. The measured height did not reach the experimental c-axis lattice parameter of bulk MnBr2, which is 6.271 Å[57]. The situation is similar for second-layer islands. The apparent heights are in the range of 4 Å- 6 Å, dependent on $V_{\mathrm{b}}$, $I_{\mathrm{t}}$, and the tip state. For an example height profile, see Figure S1 of the SI. The situation is similar to that of CoCl2, where different apparent heights ranging from 4.8 Å to 8.0 Å and a strong bias dependence have been reported [11, 10, 15].
Figure 7(b) displays a differential conductance ($\mathrm{d}I/\mathrm{d}V$) spectrum acquired from a defect-free region more than 10 nm from the edges of single-layer MnBr2 (blue data). The differential conductance is, in the first approximation, proportional to the local density of states of the sample. The conduction band edge (CBE) is located at $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and the valence band edge (VBE) at $V_{\mathrm{b}}=(1.9\pm 0.1)$ V. The resulting total gap of $(4.4\pm 0.2)$ eV makes MnBr2 an insulator, consistent with previous DFT calculations that found a band gap of 4.21 eV [24] or 3.95 eV [25].
Figure 7(b) also displays the $\mathrm{d}I/\mathrm{d}V$ spectrum of bilayer MnBr2 (red data). The CBE is located at $V_{\mathrm{b}}=(-3.0\pm 0.1)$ V and the VBE at $V_{\mathrm{b}}=(2.1\pm 0.1)$ V. The total resulting gap of $(5.1\pm 0.2)$ eV is larger than the band gap of the single-layer, contrary to the case of MoS2 [62]. The larger band gap of the bilayer is at variance with DFT calculations. They yielded a smaller band gap for the MnBr2 bulk than for the single-layer [25]. However, a similar trend has been reported for CoCl2 grown on HOPG (see Supporting Information of ref. 11), where the bilayer displays a larger band gap than the single-layer.
Evidently, the strong variation in the apparent height, as shown in Figure 7(a), is related to the electronic structure of MnBr2. In the range of the band gap between $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and $V_{\mathrm{b}}=(1.9\pm 0.1)$ V tunneling is between the tip and the corrugated moiré substrate Gr/Ir(110). The latter may be hybridized with the Br sublattice in contact with it, thus providing a mechanism for the MnBr2 layer to enter into a lateral modulation of the tunneling probability. The tunneling barrier within MnBr2 is defined by its CBE being lower than that of the vacuum barrier. This gives rise to the apparent non-zero height of MnBr2. The modulation of the barrier height by the distribution of electronic states in the CBE is consequently the second mechanism by which MnBr2 modulates the tunneling probability laterally. Both mechanisms are primarily electronic in origin and are not related to the variation of the height of the nuclei in the MnBr2 layer with lateral position. This is supported by the fact that the apparent corrugation measured in the supermoiré with 0.7 Å  is much larger than the corrugation of the Gr/Ir(110) moiré with 0.2 Å [see Figure 6(c)].
When $V_{\mathrm{b}}$ decreases below 2.5-2.5 V, a new tunneling channel into the MnBr2 valence band opens. However, it is far below the Fermi level of the sample, such that its effect on the overall tunneling probability can be neglected compared to tunneling out of the Gr states near the sample Fermi level. When $V_{\mathrm{b}}$ increases above $1.9$ V, tunneling into the MnBr2 valence band sets in, thus giving rise to an increased apparent height. The decrease in apparent height beyond $V_{\mathrm{b}}=3.2$ V indicates that above 3.2 eV the density of states in MnBr2 decreases again, as also evident in the DFT calculations of ref. 25.
III Conclusions
MnBr2 growth on Gr/Ir(110) exhibits the hallmark of van der Waals epitaxy, with substrate alignment improving at higher growth temperatures. At 400 K growth, MnBr2 is nearly epitaxially aligned with Gr, with densely packed rows rotated by 30 and exhibits a hexagonal lattice (3.90 ± 0.01 Å), consistent with bulk 1T-MnBr2. Morphology evolves from fractal to compact dendritic-skeletal islands as temperature increases. Nucleation is homogeneous at low temperatures but becomes heterogeneous at defects and step edges at room temperature and above. Second-layer nucleation is more probable than on bare Gr, suggesting stronger MnBr2-MnBr2 interaction than MnBr2-Gr, while step-edge barriers are insignificant. Upon annealing, sublimation begins at 570 K and MnBr2 is fully desorbed by 670 K.
A complex super-moiré arises from the interplay of MnBr2, Gr, and Ir(110). The strongest Fourier component corresponds to a “virtual” moiré between MnBr2 and Ir(110), despite the lattices not being in contact. This puzzle is resolved by recognizing that the dense-row periodicity of Ir(110) strongly modulates the Gr charge density, which couples with the Br sublattice of MnBr2. The resulting 13.5 nm periodicity originates from the beating between stripes of Gr/Ir(110) (3.3 nm) and MnBr2/Ir(110) (2.7 nm), while contributions from hexagonal MnBr2/Gr moiré (2.5 nm) is also required.
STS reveals a band gap of single-layer MnBr2 of approximately 4.4 eV with the valence band at $\approx$ -2.5 eV and conduction band at $\approx$ +1.9 eV. The apparent STM height is 5.95 Å , almost reaching the bulk interlayer spacing, when tunneling through the conduction band, whereas it remains rather constant at the low value of 4 Å. This behavior, along with the spatial modulation of the tunneling barrier height and hybridization of the Br sublattice with Gr/Ir(110), demonstrates that the super-moiré is predominantly electronic rather than structural in origin. The large apparent corrugation compared to that of Gr/Ir(110) further supports this conclusion.
Overall, this comprehensive characterization of an example 2D TMDH will be useful for understanding the growth, structure, and super-moirés of other 2D materials in this class. The first characterization of a super-moiré composed of lattices with different symmetries broadened our understanding of how super-moirés can emerge. In addition, virtual "moirés" of lattices that are not in contact with each other may be of primary relevance in the formation of super-moirés. Since single-layer MnBr2 is expected to host magnetic order and polarons, it will be interesting to explore whether any of these are affected by the electronic super-moiré of MnBr2/Gr/Ir(110).
IV Methods
The experiments were carried out in two ultrahigh vacuum systems with base pressures below $1\times 10^{-10}$ mbar. Each system was equipped with standard MBE growth facilities, LEED or microchannel plate LEED (MCP-LEED), and STM operated at temperatures of 300 K, 77 K, and 1.7 K.
Ir(110) is cleaned and prepared in its unreconstructed state by cycles of 4.5 keV Xe+ ion sputtering, flash annealing to 1510 K, and subsequent cooling to 400 K in $1\times 10^{-7}$ mbar oxygen pressure.
Single-crystal Gr was grown by heating cleaned unreconstructed Ir(110) to 1510 K and exposing it to $3\times 10^{-7}$ mbar ethylene for 240 s [56]. The quality of the as-grown Gr/Ir(110) is checked by LEED and STM.
MnBr2 is grown by sublimation of MnBr2 molecules from MnBr2 powder in a Knudsen cell heated to 670 K. The Knudsen cell was placed 8 cm from the sample, producing a deposition rate of $3\times 10^{-4}$ ML/s. Here, a monolayer (ML) corresponds to complete surface coverage by a single layer of MnBr2, equivalent to $7.6\times 10^{18}$ MnBr2 molecules m-2 s-1.
LEED patterns were acquired at 300 K using electron energies ranging from 90 to 150 eV. For the MCP-LEED measurements, distortions in the reciprocal space due to the flat microchannel plate geometry were corrected using a Python script.
STM was performed at 300 K in a variable temperature and at 77 K and 1.7 K in a bath cryostat STM system. Constant-current topographies were recorded with sample bias $V_{\mathrm{b}}$ and tunneling current $I_{\mathrm{t}}$, as detailed in the respective figure captions. The STM images were analyzed and processed (plane subtraction, contrast correction) using the WSxM software [63]. STS spectra were obtained at 1.7 K with stabilization bias $V_{st}$ and stabilization current $I_{st}$ using the standard lock-in technique with modulation frequency $f_{mod}$ and modulation voltage $V_{mod}$, as specified in the captions.
Our ab initio density functional theory (DFT) [64, 65] calculations were carried out using the projector augmented wave (PAW) method [66] as implemented in the Vienna Ab initio Simulation Package (VASP) [67, 68, 69]. To account for the van der Waals interactions in the Gr/Ir(110) system, we employed the non-local vdW-DF2 correlation energy functional [70] combined with a re-optimized B86b exchange functional [71, 72]. Gr on Ir(110) was modeled using a slab comprising three Ir layers and a vacuum region of $\approx 21$ Å amounting to 350 C and 264 Ir atoms. The ground-state geometry and electronic structure were determined using a plane-wave kinetic energy cutoff of 500 eV and a threshold for Hellmann-Feynman forces of $\approx 0.005$ eV/Å. Structural relaxation was performed using $\Gamma$-point sampling of the Brillouin zone, whereas the density of states was obtained using a $\times2$ $k$-point mesh.
V acknowledgments
Funding from the Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01, C01, and B06) is acknowledged. J.F. acknowledges financial support from the DFG through project FI 2624/1-1 (project no. 462692705) within SPP 2137. Financial support by the DFG priority program SPP2244 is acknowledged W.J. (project no. 535290457). The authors acknowledge the computing time granted by the JARA Vergabegremium and provided on the JARA Partition part of the supercomputer JURECA at Forschungszentrum Jülich.
Conflict of Interest
The authors declare no conflicts of interest.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author, upon reasonable request.